Let X = {1,2,3,4,5,6,7} and R = {x,y/x–y is divisible by 3} in x. Show that R is an equivalence relation. 


1
Expert's answer
2021-09-21T12:25:46-0400

Let "A = \{1, 2, 3, 4, 5, 6, 7\}" and "R = \{(x, y) | x –y \text{ is divisible by }3\}"

"4-1=3" is divisible by 3.

"5-2=3" is divisible by 3.

"6-3=3" is divisible by 3.

"7-4=3" is divisible by 3.

And vice versa.

"1-4=-3" is divisible by 3.

"2-5=-3" is divisible by 3.

"3-6=-3" is divisible by 3.

"4-7=-3" is divisible by 3.

Also,

"1-1=0" is divisible by 3.

"2-2=0" is divisible by 3.

"3-3=0" is divisible by 3.

"4-4=0" is divisible by 3.

"5-5=0" is divisible by 3.

"6-6=0" is divisible by 3.

"7-7=0" is divisible by 3.

"R=\{(4,1),(5,2),(6,3),(7,4),(1,4),(2,5),(3,6),(4,7),
\\ (1,1),(2,2),(3,3),(4,4),(5,5),(6,6),(7,7)\}"

Reflexive:

Clearly, "\{(1,1),(2,2),(3,3),(4,4),(5,5),(6,6),(7,7)\}"

So, "\{(a,a)\in R, \forall a\in A\}"

Hence, it is reflexive.


Symmetric:

Clearly, "\{(1,4),(4,1),(2,5),(5,2),(3,6),(6,3),(4,7),(7,4)\}"

So, "\{(a,b)\in R \Rightarrow (b,a)\in R, \forall a\in A\}"

Hence, it is symmetric.


Transitive:

Clearly,

"\{(1,4),(4,1),(1,1),(2,5),(5,2),(2,2),(3,6),(6,3),(3,3),(4,7),(7,4),(4,4)\}"

So, "\{(a,b)\in R, (b,c)\in R\Rightarrow (a,c)\in R,\forall a\in A\}"

Hence, it is transitive.


Thus, the given relation is an equivalence relation.